OpenStudy (anonymous):

place each column that the the number belongs in.. 2 and 3 8ths......Natural, whole integer rational irrational real........can someone please help me!:)

jimthompson5910 (jim_thompson5910):

2 and 3/8 is a rational number since... 2 and 3/8 = 2+3/8 2 and 3/8 = 16/8+3/8 2 and 3/8 = 19/8 which is a ratio of two integers It's also a real number

OpenStudy (anonymous):

ok i get that its a real number and a rational number. but where did you get the 19 and 16 from?

OpenStudy (anonymous):

ok nevermind i had got a little lost

jimthompson5910 (jim_thompson5910):

so you see how I got all that?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np